The speed of one aircraft to that of another is 2 to 5. If the slower aircraft has a speed of 300 knots, what is the speed of the faster aircraft?: The speed of one aircraft to that of another is 2 to 5. If the slower aircraft has a speed of 300 knots, what is the speed of the faster aircraft?
Answer by jojo14344(879)

See the Ratio of the slower aircraft=sa to the faster aircraft=fa:
sa/fa=2/5
Then:2*300knots=600knots, smaller aircraft
5*300=1500 knots, faster aitrcraft
:
2/5=600/1500
0.40=0.40
Thank you,
Jojo